# Decoding Generalization from Memorization in Deep Neural Networks

### v1

@@ A.10 SGD vs Adam
Figure 24: Class-wise number of PCA components of the corrupted training subspace over the layers of generalized MLP network with various corruption degrees. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with $SGD$ and $Adam$ optimizer as noted.
A.11 Regularization experiments
We ran some preliminary experiments with Dropout as a regularizer, which we describe below. The CNN model and AlexNet model were slightly modified to add regularization in the form of dropout layers. A dropout layer with dropout probability of 0.2 was used after every fully connected layer in the CNN. This model was trained on MNIST, FashionMNIST, and CIFAR-10. For the AlexNet model, likewise, dropout layers with Dropout probability of 0.5 was added after every fully connected layer in the network and the model was trained on CIFAR-100.
The results for CNN trained on MNIST, Fashion-MNIST with and without Dropout are shown in Figure 25, 29, 33 and their respective number of PCA components are shown in Figure 26, 30, 34.
The results for CNN trained on CIFAR-10 and AlexNet trained on CIFAR100, with and without drop out, are shown in Figure 27, 31, 35 and their respective number of PCA components are shown in Figure 28, 32, 36.
The results appeared somewhat inconclusive. In particular, the models we trained with Dropout did not have significantly better test accuracy than the corresponding models trained without Dropout. In some cases, the models with Dropout did not converge to high training accuracy in spite of training for 500 epochs. We conclude that this calls for more detailed study with more careful hyperparameter tuning.
Figure 25: MASC accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree, for multiple CNN models trained with and without dropout. Rows corresponds to plots with the same corruption degree & the columns correspond to the models, as noted. Training accuracy (dashed line) & testing accuracy (dotted line) of the model is shown. FC corresponds to fully connected layer with $ReLU$ activation.
Figure 26: Class-wise number of PCA components of the corrupted training subspace over the layers of CNN networks trained with and without dropout, for various corruption degrees. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted.
Figure 27: MASC accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree, for CNN and AlexNet models trained with and without dropout. Rows corresponds to plots with the same corruption degree & the columns correspond to the models as noted. Training accuracy (dashed line) & testing accuracy (dotted line) of the model is shown. FC corresponds to fully connected layer with $ReLU$ activation hereas Flat corresponds to flatten layer without $ReLU$ activation.
Figure 28: Class-wise number of PCA components of the corrupted training subspace over the layers of CNN and AlexNet networks trained with and without dropout, for various corruption degrees. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted.
Figure 29: MASC accuracy over the layers of the CNN network, trained with and without dropout, when the data set is projected onto subspace corresponding to true training labels. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Training and testing accuracy of the model is shown. FC corresponds to fully connected layer with $ReLU$ activation whereas Flat corresponds to flatten layer without $ReLU$ activation.
Figure 30: Class-wise number of PCA components of the subspace corresponding to true training labels over the layers of CNN networks trained with and without drop out and various corruption degrees. Although it is not mentioned in the legend, all the 100 classes of CIFAR-100 are plotted. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted.
Figure 31: MASC accuracy over the layers of the CNN and AlexNet, trained with and without dropout, when the data set is projected onto subspace corresponding to true training labels. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Training and testing accuracy of the model is shown. FC corresponds to fully connected layer with $ReLU$ activation whereas Flat corresponds to flatten layer without $ReLU$ activation.
Figure 32: Class-wise number of PCA components of the subspace corresponding to true training labels over the layers of CNN and AlexNet models trained with and without drop out and various corruption degrees. Although it is not mentioned in the legend, all the 100 classes of CIFAR-100 are plotted. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted.
Figure 33: MASC accuracy over the layers of the generalized CNN network, trained with and without drop out, when the data set is projected onto corrupted training subspaces with the indicated corruption degree. Rows corresponds to plots which have the same corruption degree & the columns correspond to the generalized models as noted. Training & testing accuracy of the generalized model is shown. FC corresponds to fully connected layer with $ReLU$ activation.
Figure 34: Class-wise number of PCA components of the corrupted training subspace over the layers of generalized CNN networks with various corruption degrees. Although it is not mentioned in the legend, all the 100 classes of CIFAR-100 are plotted. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with and without drop out as noted.
Figure 35: MASC accuracy over the layers of the generalized CNN and AlexNet network, when the data set is projected onto corrupted training subspaces with the indicated corruption degree. Rows corresponds to plots which have the same corruption degree & the columns correspond to the generalized models trained with and without drop out as noted. Training & testing accuracy of the generalized model is shown. FC corresponds to fully connected layer with $ReLU$ activation.
Figure 36: Class-wise number of PCA components of the corrupted training subspace over the layers of generalized CNN and AlexNet networks with various corruption degrees. Although it is not mentioned in the legend, all the 100 classes of CIFAR-100 are plotted. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with and without drop out as noted.
A.12 Tiny ImageNet : Accuracy on 0.9 and 0.99 variance captured.
MASC accuracies for the three different experiments for AlexNet trained on Tiny ImageNet with 99% variance captured and 90% variance captured is shown in Figure 37, 38, and 39. All the plots show that MASC accuracy on corrupted training, MASC accuracy on original training and MASC accuracy on testing is dependant on the percentage of variance captured by the PCA components. It can be observed by the decrease in training and testing accuracy from 99% variance to 90% variance captured plots. Although we have done this experiment on only one model and dataset, further investigation involving the the impact of number of PCA components to the MASC accuracies will be instructive.
Figure 37: MASC accuracy over the layers of the AlexNet model trained on Tiny ImageNet when the data is projected onto 99% and 90% variance explained corrupted training subspaces with the indicated corruption degree, for multiple models/datasets. Rows corresponds to plots with the same corruption degree & the columns correspond to the models, as noted. Training accuracy (dashed line) & testing accuracy (dotted line) of the model is shown. FC corresponds to fully connected layer with $ReLU$ activation whereas Flat corresponds to flatten layer without $ReLU$ activation.
Figure 38: MASC accuracy over the layers of the AlexNet models trained on Tiny ImageNet when the data set is projected onto 99.9%, 99% and 90% variance explained subspace corresponding to true training labels. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models as noted. Training and testing accuracy of the model is shown. FC corresponds to fully connected layer with $ReLU$ activation whereas Flat corresponds to flatten layer without $ReLU$ activation.
Figure 39: MASC accuracy over the layers of the generalized AlexNet model trained on Tiny ImageNet when the data set is projected onto 99% and 90% variance explained training subspaces with the indicated corruption degrees. Rows corresponds to plots which have the same corruption degree & the columns correspond to the generalized models as noted. Training & testing accuracy of the generalized model is shown. FC corresponds to fully connected layer with $ReLU$ activation whereas Flat corresponds to flatten layer without $ReLU$ activation.

### v2

@@ Appendix H Investigating latent memorization capabilities of uncorrupted models
Figure 41: Class-wise number of PCA components of the corrupted training subspace over the layers of generalized MLP network with various corruption degrees. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with SGDSGD and AdamAdam optimizer as noted.
Appendix L Retraining experiment results on AlexNet models and additional analysis results.
Retraining the model using MASC experiment results on AlexNet models i.e, AlexNet-CIFAR-100 and AlexNet-Tiny ImageNet for different corruption degrees are shown in this section. Test accuracies averaged over three runs on the 80% test dataset is plotted for different AlexNet-dataset pairs for various corruption degrees and for various models/MASC classifiers is shown in Figure 42.
Decrease in fraction of incorrectly labeled data points resulting from relabeling the training set expressed as a fraction of size incorrectly labeled points in the existing corrupted data, using the corresponding best-layer MASC classifier (corrupted subspace and true subspace) is shown in Figure 43 for all models-dataset pairs and corruption degree 20%, 40%, 60%, & 80%. The average values for reference are available in Table 9. The percentage of incorrect labels that are correctly relabeled using MASC (corrupted subspace and true subspace) for all models-dataset and corruption degrees 20%, 40%, 60%, 80% is shown in Figure 44. The average values for reference are available in Table 10. The percentage of incorrect labels that are correctly relabeled using MASC (corrupted subspace and true subspace) for all models-dataset and corruption degrees 20%, 40%, 60%, 80% is shown in Figure 45. The average values for reference are available in Table 11.
Figure 42: Test accuracies averaged over three runs on the 80% test dataset is plotted for different AlexNet-dataset pairs for various corruption degrees and for various models/MASC classifiers. Model before retraining corresponds to the existing memorized model. Model after retraining on MASC corrupted subspace corresponds to model trained with training dataset relabeled using MASC corrupted subspace predictions on the best layer. Model after retraining on MASC true subspace corresponds to model trained with training dataset relabeled using MASC subspaces corresponding to true label predictions on the best layer. MASC test accuracy on the best layers for corrupted and true label subspaces on existing corrupted models (before retraining) are shown for comparison. The best layer was identified using the validation set which was carved out of the test set for this experiment. Error bar represents the range on three different runs.
Figure 43: Decrease in fraction of incorrectly labeled data points resulting from relabeling the training set expressed as a fraction of size incorrectly labeled points in the existing corrupted data, using the corresponding best-layer MASC classifier. That is if X is the number of data points incorrectly labeled in corrupted dataset and Y is the number of data points incorrectly labeled in relabeled dataset then each of these plots refer to (X-Y)/X. Rows corresponds to results with the same corruption degree. The first column correspond to results with MLP and CNN models and second column with AlexNet model, as noted. The fractions are averaged over three runs. Error bar represents the range on three different runs.
Table 9: The average decrease in fraction of incorrectly labeled data points resulting from relabeling the training set expressed as a fraction of size incorrectly labeled points in the existing corrupted data, using the corresponding best-layer MASC classifier. The values are rounded to the second decimal place.
Subspace Corrupted labels True labels
Corruption degree 20% 40% 60% 80% 20% 40% 60% 80%
MLP-MNIST 76.93% 38.63% 19.90% 7.75% 97.83% 99.66% 99.93% 99.95%
MLP-CIFAR-10 3.45% 2.38% 1.10% 0.22% 95.49% 97.91% 98.75% 99.14%
CNN-MNIST 88.15% 89.19% 86.44% 71.12% 92.42% 96.46% 96.93% 97.12%
CNN-Fashion-MNIST 40.69% 64.96% 69.38% 62.61% 35.96% 66.08% 79.26% 80.59%
CNN-CIFAR-10 6.25% 3.30% 9.18% 3.79% 16.44% 40.22% 81.75% 80.09%
AlexNet-CIFAR-100 -0.09% 0.26% -0.20% -0.32% 99.61% 98.96% 90.20% 95.13%
AlexNet-Tiny ImageNet -92.46% -26.50% -3.95% -11.55% -91.32% -20.24% 2.89% -12.38%
Figure 44: The percentage of incorrect labels that are correctly relabeled using MASC (corrupted subspace and true subspace) for all the models and for corruption degrees 20%, 40%, 60%, 80%. Rows corresponds to results with the same corruption degree. The first column correspond to results with MLP and CNN models and second column with AlexNet model, as noted. The percentage are averaged over three runs. Error bar represents the range on three different runs.
Table 10: The average percentage of incorrect labels that are correctly relabeled using MASC (corrupted subspace and true subspace) for all models and for corruption degrees 20%, 40%, 60%, 80%. The values are rounded to the second decimal place.
Subspace Corrupted labels True labels
Corruption degree 20% 40% 60% 80% 20% 40% 60% 80%
MLP-MNIST 77.62% 38.65% 19.90% 7.75% 99.23% 99.80% 99.95% 99.96%
MLP-CIFAR-10 5.59% 2.82% 1.20% 0.24% 99.08% 99.12% 99.23% 99.37%
CNN-MNIST 93.71% 91.34% 87.64% 72.08% 98.27% 98.54% 98.26% 97.88%
CNN-Fashion-MNIST 81.42% 79.18% 76.16% 66.23% 83.82% 86.80% 88.76% 86.08%
CNN-CIFAR-10 17.70% 11.54% 12.30% 5.51% 41.98% 55.63% 89.85% 85.41%
AlexNet-CIFAR-100 0.15% 0.43% 0.19% 0.06% 99.84% 99.32% 91.06% 95.34%
AlexNet-Tiny ImageNet 5.94% 3.96% 0.51% 0.41% 25.01% 24.43% 18.77% 3.26%
Figure 45: The percentage of correct labels that are incorrectly relabeled using MASC (corrupted subspace and true subspace) for all the models and for corruption degrees 20%, 40%, 60%, 80%. Rows corresponds to results with the same corruption degree. The first column correspond to results with MLP and CNN models and second column with AlexNet model, as noted.The percentage are averaged over three runs. Error bar represents the range on three different runs.
Table 11: The average percentage of correct labels that are incorrectly relabeled using MASC (corrupted subspace and true subspace) for all models and for corruption degrees 20%, 40%, 60%, 80%. The values are rounded to the second decimal place.
Subspace Corrupted labels True labels
Corruption degree 20% 40% 60% 80% 20% 40% 60% 80%
MLP-MNIST 0.15% 0.01% 0% 0% 0.31% 0.08% 0.03% 0.03%
MLP-CIFAR-10 0.47% 0.25% 0.13% 0.06% 0.79% 0.69% 0.58% 0.59%
CNN-MNIST 1.22% 1.20% 1.40% 2.49% 1.28% 1.17% 1.55% 1.97%
CNN-Fashion-MNIST 8.86% 8.06% 8.04% 9.30% 10.42% 11.76% 11.26% 14.12%
CNN-CIFAR-10 2.52% 4.67% 3.69% 4.48% 5.62% 8.73% 9.60% 13.82%
AlexNet-CIFAR-100 0.06% 0.11% 0.57% 1.44% 0.06% 0.23% 1.26% 0.81%
AlexNet-Tiny ImageNet 24.46% 20.10% 6.60% 46.97% 28.92% 29.48% 23.51% 61.39%
Appendix M Retraining experiment: epoch-wise results
To study the dynamics of accuracy during retraining, unencumbered by the early stopping criterion, we also performed a similar experiment without using early stopping, for 10 epochs. The results for model before training, model after retraining on MASC corrupted subspace, and model after retraining on MASC original subspace for all model-dataset pair and for various corruption degrees are shown in 46. The following steps are performed to retrain the model for corrupted and original subspaces independently.
  1. 1.
    Test dataset is split into 80%-20%.
  2. 2.
    The model’s best layer is identified using MASC accuracy on the validation set (carved out from 20% of the test dataset).
  3. 3.
    The corrupted training dataset was relabeled using the best layer MASC predictions.
  4. 4.
    The existing corrupted model was loaded for retraining purpose. Using the relabeled corrupted training dataset, the model was further trained for 10 epochs.
  5. 5.
    The remaining 80% test dataset was used to calculate generalization to true labels.
Figure 46: Test accuracy on 80% dataset is plotted for different model-dataset pairs for various corruption degrees over 10 epochs. The experiment was performed without using early stopping. Model before retraining corresponds to the existing memorized model. Model after retraining on MASC corrupted subspace corresponds to model trained with training dataset relabeled using MASC corrupted subspace predictions on the best layer. Model after retraining on MASC true subspace corresponds to model trained with training dataset relabeled using MASC subspaces corresponding to true label predictions on the best layer. Error bar represents the range on three different runs.